OpenStudy (anonymous):

Which expression is equivalent to 11(17 – x)? A. 11 • 17 – 11x B. 11 • 17 – x C. 17 – 11x D. 11 + 17 – x

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

help please

OpenStudy (igreen):

11(17 - x) is the same thing as multiplying 11 to everything in the parenthesis, which means you multiply 11 to 17 and -x

OpenStudy (anonymous):

i need more help after this please too

OpenStudy (anonymous):

so........B. then?

OpenStudy (anonymous):

???????????????????????

OpenStudy (igreen):

No, close..

OpenStudy (anonymous):

A?

OpenStudy (igreen):

Yep!

OpenStudy (anonymous):

i more please hold on please

OpenStudy (igreen):

11 * -x = -11x

OpenStudy (igreen):

Okay.

OpenStudy (anonymous):

Combine like terms. 8 + 6x – 11x + 12 A. 20 + 5x B. 20 – 17x C. 14x + 1x D. 20 – 5x

OpenStudy (anonymous):

leave then come back in to see it better sorry:)

OpenStudy (anonymous):

is it D.???

OpenStudy (anonymous):

20-5x?

OpenStudy (igreen):

Hold on.

OpenStudy (anonymous):

k

OpenStudy (igreen):

6 - 11 = ? 8 + 12 = ?

OpenStudy (anonymous):

17 20

OpenStudy (igreen):

That's 6 - 11, not 6 + 11.

OpenStudy (anonymous):

-5

OpenStudy (igreen):

Yep, so that gives us -5x and 20. Or 20 - 5x.

OpenStudy (anonymous):

ok mi have another please

OpenStudy (igreen):

Okay, sure.

OpenStudy (anonymous):

Evaluate the expression by using the given value of the variables. y ÷ 2 • 4 – 10 ÷ x + 6 where x = 5 and y = 16 A. 36 B. 3 C. –3 D. 21

OpenStudy (anonymous):

is it a.

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

thanks bye:)

OpenStudy (igreen):

??

OpenStudy (igreen):

\(y \div 2 \cdot 4 – 10 \div x + 6\) x = 5 y = 16 \(16 \div 2 \cdot 4 – 10 \div 5 + 6\) \(8 \cdot 4 – 10 \div 5 + 6\) \(32 – 10 \div 5 + 6\) \(32 – 2 + 6\) \(30 + 6\) \(36~\huge\color{lime} \checkmark\)